If a function f:[ a,b] to R has finitely many points of discontinuity in [ a,b] , then f is integrable on [ a,b].


True or false with full explanation

1
Expert's answer
2021-11-17T03:47:34-0500

Solution:

True.

Proof:

Suppose "c \in(a, b)" . Let "\varepsilon>0" . Since f is bounded, there exists M>0 such that "|f(x)| \leq M" for all "x \in[a, b]" . Let "\delta=\varepsilon / 12 M" . Since f is continuous on"[a, c-\delta]" and "[c+\delta, b]" , f is Riemann-integrable on "[a, c-\delta]" and "[c+\delta, b]" , so there exists partitions "P_{1}=\left\{a=x_{0}<\cdots<x_{n}=c-\delta\right\} of [a, c-\delta]" and "P_{2}=\left\{c+\delta=y_{0}<\cdots<y_{m}=c-\delta\right\}" of "[c+\delta, b]" such that

"U\left(P_{1}, f\right)-L\left(P_{1}, f\right)<\frac{\varepsilon}{3} \quad \text { and } \quad U\left(P_{2}, f\right)-L\left(P_{2}, f\right)<\frac{\varepsilon}{3}"

Consider the partition of [a, b] given by "P=P_{1} \cup P_{2}" . Then

"\begin{aligned}

U(P, f) &=\sum_{i=1}^{n} M_{i} \Delta x_{i}+2 \delta \sup _{x \in[c-\delta, c+\delta]} f(x)+\sum_{j=1}^{m} M_{j} \Delta y_{j} \leq U\left(P_{1}, f\right)+2 M \delta+U\left(P_{2}, f\right) \\

L(P, f) &=\sum_{i=1}^{n} m_{i} \Delta x_{i}+2 \delta \inf _{x \in[c-\delta, c+\delta]} f(x)+\sum_{j=1}^{m} m_{j} \Delta y_{j} \geq L\left(P_{1}, f\right)-2 M \delta+L\left(P_{2}, f\right)

\end{aligned}"

and so

"\begin{aligned}

U(P, f)-L(P, f) & \leq\left[U\left(P_{1}, f\right)-L\left(P_{1}, f\right)\right]+4 M \delta+\left[U\left(P_{2}, f\right)-L\left(P_{2}, f\right)\right] \\

&<\frac{\varepsilon}{3}+4 M \times \frac{\varepsilon}{12 M}+\frac{\varepsilon}{3}=\varepsilon

\end{aligned}"

Hence f is Riemann integrable.